OpenStudy (raheelafzaal):

find the greatest number which divides 742 and 1162 leaving 7 as reminder in each case?

OpenStudy (anonymous):

Answer: This is a slight twist to the normal find the GCF of two numbers. In this case as a remainder of 7 is required, subtracting 7 from each number and then finding the GCF of the resulting numbers will solve the problem: 742 - 7 = 735 1162 - 7 = 1155 GCF of 1155 and 735 (using Euclid's method): 1155 / 735 = 1 r 420 735 / 420 = 1 r 315 420 / 315 = 1 r 105 315 / 105 = 3 r 0 GCF of 735 & 1155

OpenStudy (raheelafzaal):

i have 1 more question

OpenStudy (anonymous):

yes?